# Multi-Level Embedding and Alignment Network with Consistency and Invariance Learning for Cross-View Geo-Localization

### v1

@@ Abstract
Cross-View Geo-Localization (CVGL) involves determining the localization of drone images by retrieving the most similar GPS-tagged satellite images. However, the imaging gaps between platforms are often significant and the variations in viewpoints are substantial, which limits the ability of existing methods to effectively associate cross-view features and extract consistent and invariant characteristics. Moreover, existing methods often overlook the problem of increased computational and storage requirements when improving model performance. To handle these limitations, we propose a lightweight enhanced alignment network, called the Multi-Level Embedding and Alignment Network (MEAN). The MEAN network uses a progressive multi-level enhancement strategy, global-to-local associations, and cross-domain alignment, enabling feature communication across levels. This allows MEAN to effectively connect features at different levels and learn robust cross-view consistent mappings and modality-invariant features. Moreover, MEAN adopts a shallow backbone network combined with a lightweight branch design, effectively reducing parameter count and computational complexity. Experimental results on the University-1652 and SUES-200 datasets demonstrate that MEAN reduces parameter count by 62.17% and computational complexity by 70.99% compared to state-of-the-art models, while maintaining competitive or even superior performance. The codes will be released soon.
@@
Cross-view geo-localization (CVGL) has received extensive attention in autonomous vehicles, aerial photography, and autonomous navigation [1, 2, 3]. CVGL is usually recognized as an image retrieval task on the heterogeneous platform, enabling one to accurately determine the geo-location of a query image by matching it with several reference images taken from the varying views of different platforms. It has been more common in earlier applications to achieve geo-localization by matching the ground panoramic images with satellite images [4, 5, 6]. Recently, drone aerial photography has further extended its applications [7, 8, 9], including drone target localization and drone navigation. In these cases, the images taken by drones can be matched with satellite images of the same geographic area, which contain precise longitude and latitude coordinates, achieving indirect localization of drones. This realizes drone navigation, replacing the need for global navigation satellite system (GNSS) equipment onboard [10]. However, CVGL encounters substantial challenges due to appearance variations and spatial misalignments caused by scale variations, imaging gaps, and spatial transformations in the aerial photography scene.
@@
Prior research focused on improving feature representation capabilities, initially using global features to encompass overall structural information [11, 12] or local features to detail fine-grained aspects [13, 14]. However, global features lack the ability to represent detailed information, whereas local features are sensitive to variations in viewpoint or scale. Consequently, some researches have developed the joint global-local feature representations to overcome their respective inherent limitations, incorporating more powerful feature extraction architectures. For example, the Transformer [15, 16] is used to model long-range dependencies and complex contextual semantics at both the global and local levels. Although these architectures possess strong feature extraction capabilities to enhance overall feature representation and mitigate the discrepancies caused by appearance variations, the disparity in feature space distribution due to spatial misalignment makes it difficult to obtain efficient alignment of cross-view features, namely feature consistency. Moreover, the excessive emphasis on contextual information can induce noise interference [17], limiting feature consistency and matching accuracy in the complex retrieval tasks. Furthermore, increasing architecture complexity results in a high parameter count, significantly increasing the computational costs of deployment in environments with limited airborne resources [18].
@@
Attempts to extract the cross-view consistent features provide the possibilities to accurately align and associate the same target from different viewpoints and transformed spatial [19]. Several methods [12, 20, 21] have been proposed to address inconsistencies in the feature distribution. These methods combine dense partition learning [9, 22] and feature alignment strategies [23, 19] to capture deep structural relationships of cross-view features within a shared embedding space. To further improve the discriminability of features in the cross-view mappings of these methods, contrastive loss [5] and triplet loss [24] have been used to achieve a more precise differentiation between positive and negative samples within the embedding space.
@@
Although these methods have made significant progress in addressing feature distribution inconsistencies through dense partition learning and feature alignment, they often overemphasize fine-grained feature alignment. In this case, it relied heavily on the specific details of the viewpoints or scales,making it difficult to effectively associate features across different levels. In scenarios characterized by substantial viewpoint differences or large-scale spatial variations, consistency mappings may not capture the critical cross-view invariant that differs from these environmentally sensitive features, i.e., the invariance of features[25]. This limited their robustness and uniformity, ultimately undermining the generalization capability.
@@
To address those issues, we propose a novel enhanced alignment network for the cross-view geo-localization tasks, called multi-level embedding alignment network (MEAN). In the proposed MEAN, ConvNeXt-Tiny, the smallest network in the ConvNeXt family[26] is used as a backbone to extract coarse-grained features. These features are then processed separately by the three designed branches. The progressive diversification embedding branch proposed in the MEAN adopts a stepwise expansion strategy to generate a rich set of embedding features, allowing the model to learn diverse feature representations. Furthermore, it uses contrastive loss to enhance the consistency and discriminability of the shared embeddings. To better coordinate feature representation capabilities, the other branch, named the global expansion embedding branch, jointly optimizes global and fine-grained feature representations to realize the global-to-local associations. Additionally, the multi-level feature fusion and adaptive calibration strategies with multi-level constraint achieve precise and robust feature alignment in the embedding space in the cross-domain enhanced alignment branch. Therefore, the proposed MEAN not only enhances feature representation, but also deeply explores the invariant modes and potential commonalities within the features. As shown in Fig. 1, our method significantly achieves improvements in both matching accuracy and computational efficiency compared with existing state-of-the-art methods. Its main contributions are summarized as follows.
@@ II-A Feature Extraction and Context Enhancement
The first category of methods focused primarily on enhancing feature representation, typically achieving cross-view geo-localization through single global or local feature extraction. LCM [11] mapped drone and satellite images into a unified feature space, transforming the task into a global location classification to capture the overall structure of the image. However, the ability to learn fine-grained features is limited by such global methods. Therefore, LPN [13] introduced a local pattern partitioning method that segmented images into multiple distance-based regions using a square ring partitioning strategy. This method effectively extracts fine-grained contextual information surrounding the target. Although this local method improves the model’s sensitivity to detail, single global or local feature extraction still proves limited in handling significant viewpoint and scale variations. Moreover, relying on a single branch for processing either global or local features hinders effective integration of multi-scale and multi-level spatial information in cross-view scenarios. To address these issues, IFSs [8]proposed a multi-branch joint representation learning strategy that integrates global and local feature information through a multi-branch structure.
@@
However, these methods are mainly based on convolutional neural networks (CNNs), which are inherently limited in modeling long-range dependencies and complex contextual semantics, particularly in cross-view geo-localization tasks[22]. Transformer and ConvNeXt architectures have been gradually adopted in cross-view geo-localization due to their advantages in long-range dependency modeling and contextual information extraction. TransFG[10] employed Transformer-based feature aggregation and gradient-guided modules to effectively integrate global and local information. SRLN[27] based on Swin Transformer [28] combined direction guidance and multi-scale feature fusion strategies, effectively bridging viewpoint and scale discrepancies. MCCG [7], in contrast, introduced a multi-classifier method based on ConvNeXt to learn rich feature representations, though the emphasis on contextual information can lead to over-attention to non-essential features. To mitigate this, CCR [29] introduced counterfactual causal reasoning to strengthen the model’s attention mechanism, helping it distinguish between essential and non-essential features. MFJR[17] used a multi-level feedback joint representation learning method, incorporating an adaptive region elimination strategy to effectively suppress irrelevant information and focus on key target features.
@@
Despite the effectiveness of Transformer-based and large ConvNeXt architectures, combined with specific multi-level learning strategies, in capturing features, their high computational costs pose challenges for deployment in resource-constrained environments. Furthermore, with significant viewpoint or scale variations, maintaining robustness and consistency in cross-view matching remains difficult.
@@ II-B Feature Alignment and Discriminative Optimization
To extract consistency across varying cross-view features, researchers have sought to employ spatial alignment strategies and learn more consistent and discriminative features within a shared feature space. PCL[23] utilized Perspective Projection Transformation (PPT) to align UAV and satellite images, reducing spatial misalignment between views, and then employed a CGAN to synthesize realistic satellite image styles, thereby narrowing the viewpoint gap. However, this explicit alignment method struggles to effectively learn internal feature differences and may add unnecessary noise. Consequently, many researchers have attempted to integrate the metric loss and the contrastive loss within a shared space to guide the model in capturing these internal discrepancies. However, this method is often challenged by interference from viewpoint variations and exhibits limited discriminative power for negative samples. To address this, Sample4Geo [5] employed a hard negative sampling strategy to enhance the model’s feature discrimination and contrastive learning effectiveness. However, significant differences in feature representations across cross-views challenge the extraction and alignment of geographical features under substantial viewpoint variations. Relying solely on constraint-based alignment within a shared embedding space often proves insufficient to address these discrepancies. Therefore, CAMP[30]introduced contrastive attribute mining and position-aware partitioning strategies to align geographic features under varying viewpoints and scales. Although this method enhances local feature consistency and discrimination, it excels at extracting explicit differences across viewpoints but struggles to learn feature invariance under significant viewpoint and scale changes, thereby limiting the model’s generalization capability. DAC[19]adopted domain alignment and scene consistency constraints to achieve coarse-to-fine feature consistency, relying primarily on direct alignment without fully exploring deep invariant patterns across viewpoints. As a result, DAC faces challenges in maintaining stable feature mappings under extreme viewpoint or scale variations. Moreover, both CAMP and DAC incur high computational costs in improving alignment accuracy, impacting model efficiency and deployment flexibility.
@@ III Multi-Level Embedding and Alignment Network
The proposed MEAN network is introduced in this section, with an overview of this model illustrated in Fig.2. MEAN utilizes ConvNeXt-Tiny as the backbone to extract initial coarse-grained features. The progressive extension embedding (PEE) branch and global extension embedding (GEE) branch generate multi-embedding representations, seizing representative cross-view feature cues, and refining global spatial representations, respectively. Additionally, the cross-domain enhanced alignment (CEA) branch is integrated to exhaustively extract underlying structural patterns across different view images. During the training phase, distinct loss functions are utilized for the features extracted from each branch. In particular, a cross-joint optimization strategy is employed in the PEE and CEA branches. By incorporating multi-level constraints, MEAN promotes feature enhancement and alignment, while optimizing the adaptability of feature representation for cross-view consistency and modality invariance.
Problem Formulation: Given a cross-view geo-localization dataset, we denote the input image pairs as , where and represent the images taken from drone and satellite views, respectively. Each image is associated with a label , where denotes the total number of categories corresponding to different locations in the dataset. For example, in the University-1652 dataset, there are 701 buildings, each containing one satellite view image and multiple drone view images. For the cross-view geo-localization task, the goal is to learn a mapping function that projects images from different platforms (e.g., drone and satellite) into a shared semantic space. In this space, the features of images representing the same location should be close to each other, while those from different locations should be well separated. This method enables effective matching of images from different views of the same location by leveraging feature similarity. It can address the challenges of significant viewpoint differences and supports accurate geo-localization.
@@ III-B Progressive Extension Embedding branch
Subsequently, adaptive average pooling (AvgPool) is applied to aggregate the multi-scale information, producing more discriminative semantic embeddings. Finally, the compact feature representation is processed through the batch normalization operation , followed by a dropout operation for regularization to aviod overfitting, and then propagated to the linear layer for classification. We summarize these processing operations (,,,) as
@@ III-C Global Extension Embedding Branch
The GEE branch adopts a global context aggregation strategy, compared with the PEE branch, which progressively generates and enhances local embedded details. In the GEE branch, a mean pooling operation replaces the DEG module, processing the output features of the backbone to obtain a global feature representation. The global feature representation is then propagated to the DEC module for further refinement.
@@
where, and are height factor and width factor, respectively. This operation compresses the spatial dimensions, where . The global feature is then fed into the DEC module for further processing to obtain described as DEC module III-B. In this way, the GEE branch can aggregate global information, generating comprehensive global feature embeddings that provide the DEC module with context-aware global features. Based on this global perspective, global and local consistency is ensured in the GEE branch.
@@ III-D Cross-Domain Enhanced Alignment branch
In cross-view geo-localization tasks, images taken from different platforms often exhibit significant geometric differences, resulting in substantial variations in feature distributions and posing challenges for cross-domain matching. Feature alignment has been proven to be effective in addressing this issue[31, 32, 33]. Inspired by [19], we designed a cross-domain enhanced alignment (CEA) branch that uses multi-level fusion and adaptive calibration strategies to dynamically adjust feature consistency within a shared latent space in high-dimensional embeddings. In addition, invariance constraints are incorporated to learn the inherent structural invariances of these features. As shown in Fig.2 after the initial feature extraction by the backbone, we obtain high-dimensional features from different perspectives (drone view and satellite view ). To efficiently utilize global information and facilitate cross-view feature alignment and matching, we apply a transformation to reshape these high-dimensional features into two-dimensional matrices as the input feature of the CEA branch, where and . Break down the subsequent procedures into three modules, including domain enhancement (DE) module, feature adaptive temperature (FAT) module, and multi-scale feature fusion (MSF) module, which are described as follows.

### v2

@@ Abstract
Cross-View Geo-Localization (CVGL) involves determining the localization of aerial images by retrieving the most similar GPS-tagged satellite images. However, the imaging gaps between platforms are often significant and the variations in viewpoints are substantial, which limits the ability of existing methods to effectively associate cross-view features and extract consistent and invariant characteristics. Moreover, existing methods often overlook the problem of increased computational and storage requirements when improving model performance. To handle these limitations, we propose a lightweight enhanced alignment network, called the Multi-Level Embedding and Alignment Network (MEAN). The MEAN framework uses a progressive multi-level enhancement strategy, global-to-local associations, and cross-domain alignment, enabling feature communication across levels. This allows MEAN to effectively connect features at different levels and learn robust cross-view consistent mappings and modality-invariant features. Moreover, MEAN adopts a shallow backbone network combined with a lightweight branch design, effectively reducing parameter count and computational complexity. Experimental results on the University-1652 and SUES-200 datasets demonstrate that MEAN reduces parameter count by 62.17% and computational complexity by 70.99% compared to state-of-the-art models, while maintaining competitive or even superior performance. Our code and models will be released on https://github.com/ISChenawei/MEAN.
@@
Cross-view geo-localization (CVGL) has received extensive attention in autonomous vehicles, aerial photography, and autonomous navigation [1, 2, 3]. CVGL is usually recognized as an image retrieval task on the heterogeneous platform to accurately determine the geo-location of a query image by matching it with several reference images taken from the varying views of different platforms. The early applications were concentrated on the matching of ground panoramic images with satellite images [4, 5, 6]. Recently, drone aerial photography has further extended its applications including drone target location and drone navigation [7, 8, 9]. In these cases, the aerial images taken by drones can be matched with satellite images of the same geographic area with precise longitude and latitude coordinates, achieving indirect drone location. This realizes drone navigation and replaces the need for global navigation satellite system (GNSS) equipment onboard [10]. However, CVGL encounters substantial challenges due to appearance variations and spatial misalignments caused by imaging gaps, scale variations and spatial transformations in the aerial photography scene.
@@
Previous research for CVGL focused on improving feature representation capabilities, initially using global features to encompass overall structural information [11, 12] or local features to detail fine-grained aspects [13, 14]. However, global features lack the ability to represent detailed information, whereas local features are sensitive to variations in viewpoint or scale. Consequently, some research has developed the joint global-local feature representations to overcome their respective inherent limitations with the more powerful feature extraction architectures. For example, the Transformer [15, 16] was used to model long-range dependencies and complex contextual semantics at both the global and local levels. Although these architectures possess strong feature extraction capabilities to enhance overall feature representation and mitigate the discrepancies caused by appearance variations, the disparity in feature space distribution due to spatial misalignment makes it difficult to obtain efficient alignment of cross-view features, namely feature consistency. Moreover, the excessive emphasis on contextual information can induce noise interference [17] to affect the consistency of the characteristics and the accuracy of matching in complex retrieval tasks. Furthermore, increasing architecture complexity results in a high parameter count and significantly imposes a higher computational load on limited airborne resources [18].
@@
Attempts to extract the cross-view consistent features provided the possibilities to accurately align and associate the same target from different viewpoints and transformed spatial [19]. Several methods [12, 20, 21] have been proposed to address inconsistencies in the feature distribution. These methods combined dense partition learning [9, 22] and feature alignment strategies [23, 19] to capture deep structural relationships of cross-view features within a shared embedding space. To further improve the discriminability of features in the cross-view mappings of these methods, contrastive loss [5] and triplet loss [24] have been used to achieve a more precise differentiation between positive and negative samples within the embedding space.
@@
Although these methods have made significant progress in addressing feature distribution inconsistencies through dense partition learning and feature alignment, they often overemphasized fine-grained feature alignment. In this case, it relied heavily on the specific details of the viewpoints or scales, making it difficult to effectively associate features across different levels. In scenarios characterized by substantial viewpoint differences or large-scale spatial variations, consistency mappings may not capture the critical cross-view invariant that differs from these environmentally sensitive features, i.e., the invariance of features [25]. This limited their robustness and uniformity, ultimately undermining the generalization capability.
@@
To address those issues, we propose a novel enhanced alignment network for the cross-view geo-localization tasks, called multi-level embedding alignment network (MEAN). In the proposed MEAN, ConvNeXt-Tiny, the smallest network in the ConvNeXt family [26] is used as a backbone to extract coarse-grained features. These features are then processed separately by the three designed branches. The progressive diversification embedding branch proposed in the MEAN adopts a stepwise expansion strategy to generate a rich set of embedding features, allowing the model to learn diverse feature representations. Furthermore, it uses contrastive loss to enhance the consistency and discriminability of the shared embeddings. To better coordinate feature representation capabilities, the other branch, named the global expansion embedding branch, jointly optimizes global and fine-grained feature representations to realize the global-to-local associations. Additionally, the multi-level feature fusion and adaptive calibration strategies with multi-level constraint achieve precise and robust feature alignment in the embedding space in the cross-domain enhanced alignment branch. Therefore, the proposed MEAN not only enhances feature representation, but also deeply explores the invariant modes and potential commonalities within the features. As shown in Fig. 1, our method significantly achieves improvements in both matching accuracy and computational efficiency compared to existing state-of-the-art methods. Its main contributions are summarized as follows.
@@ II-A Feature Extraction and Context Enhancement
The first category of methods focused primarily on enhancing feature representation, typically achieving cross-view geo-localization through single global or local feature extraction. LCM [11] mapped aerial and satellite images into a unified feature space, transforming the task into a global location classification to learn the overall structure of the image. However, the ability to learn fine-grained features was limited in such global methods. Therefore, LPN [13] introduced a local pattern partitioning method that segmented images into multiple distance-based regions using a square ring partitioning strategy. This method effectively extracted fine-grained contextual information surrounding the target. Although this local method improved the model’s sensitivity to detail, it was found to be limited in handling significant viewpoint and scale variations by the extraction of the single global or local feature. Moreover, relying on a single branch for processing either global or local features hindered the effective integration of multi-scale and multi-level spatial information in cross-view scenarios. To address these issues, IFSs [8] proposed a multi-branch joint representation learning strategy that integrated global and local feature information through a multi-branch structure.
@@
However, these methods were mainly based on convolutional neural networks (CNNs), which are inherently limited in modeling long-range dependencies and complex contextual semantics, particularly in cross-view geo-localization tasks [22]. Transformer and ConvNeXt architectures have been gradually adopted in cross-view geo-localization due to their advantages in long-range dependency modeling and contextual information extraction. TransFG [10] used Transformer-based feature aggregation and gradient-guided modules to effectively integrate global and local information. SRLN [27] based on Swin Transformer [28] combined direction guidance and multi-scale feature fusion strategies, effectively bridging viewpoint and scale discrepancies. MCCG [7], on the contrary, introduced a multi-classifier method based on ConvNeXt to learn rich feature representations. However, the emphasis on contextual information can lead to over-attention to non-essential features. To mitigate this, CCR [29] introduced counterfactual causal reasoning to strengthen the model’s attention mechanism, which is useful to distinguish between essential and non-essential features. MFJR [17] used a multi-level feedback joint representation learning method, incorporating an adaptive region elimination strategy to effectively suppress irrelevant information and focus on key target features.
@@
Despite the effectiveness of Transformer-based and large ConvNeXt architectures with specific multi-level learning strategies in capturing features, their high computational costs pose challenges in resource-constrained environments. Furthermore, under significant viewpoint or scale variations, it is difficult to maintain the robustness and consistency in cross-view matching.
@@ II-B Feature Alignment and Discriminative Optimization
To achieve consistency in extracting features across different cross-views, researchers have utilized spatial alignment strategies to learn more consistent and discriminative features within a shared feature space. PCL [23] utilized Perspective Projection Transformation (PPT) to align aerial and satellite images to reduce spatial misalignment between views, and then employed a CGAN to synthesize realistic satellite image styles to narrow the imaging gap. This explicit alignment method struggled to learn internal feature differences effectively and might add unnecessary noise. Consequently, many researchers have attempted to integrate the metric loss and the contrastive loss within a shared space to identify the internal discrepancies. However, this method is often challenged by interference from viewpoint variations and exhibits limited discriminative power for negative samples. To address this, Sample4Geo [5] employed a hard negative sampling strategy to improve the model’s feature discrimination and contrastive learning effectiveness. However, significant differences in feature representations across cross-views posed challenges to the extraction and alignment of geographical features under substantial viewpoint variations. Relying solely on constraint-based alignment within a shared embedding space has been shown to be insufficient to address these discrepancies. Therefore, CAMP [30] introduced contrastive attribute mining and position-aware partitioning strategies to align geographic features under varying viewpoints and scales. Although this method has enhanced local feature consistency and discrimination, it excelled in extracting explicit differences across viewpoints but struggled to learn feature invariance under significant changes in viewpoint and scale, thereby limiting the model’s generalization capability. DAC [19] adopted domain alignment and scene consistency constraints to achieve coarse-to-fine feature consistency, relying primarily on direct alignment without fully exploring deep invariant patterns across viewpoints. As a consequence, DAC encountered challenges in maintaining stable feature mappings under extreme variations in viewpoint or scale. Furthermore, both CAMP and DAC incurred high computational costs in enhancing alignment accuracy, thereby affecting model efficiency and limiting deployment flexibility.
@@ III Multi-Level Embedding and Alignment Network
The proposed MEAN framework is introduced in this section, with an overview of this model illustrated in Fig.2. MEAN utilizes ConvNeXt-Tiny as the backbone to extract initial coarse-grained features. The progressive extension embedding (PEE) branch and global extension embedding (GEE) branch generate multi-embedding representations, seizing representative cross-view feature cues, and refining global spatial representations, respectively. Additionally, the cross-domain enhanced alignment (CEA) branch is integrated to exhaustively extract underlying structural patterns across different view images. During the training phase, distinct loss functions are utilized for the features extracted from each branch. In particular, a cross-joint optimization strategy is employed in the PEE and CEA branches. By incorporating multi-level constraints, MEAN promotes feature enhancement and alignment, while optimizing the adaptability of feature representation for cross-view consistency and modality invariance.
Problem Formulation: Given a cross-view geo-localization dataset, we denote the input image pairs as , where and represent the images taken from drone and satellite views, respectively. Each image is associated with a label , where denotes the total number of categories corresponding to different locations in the dataset. For example, in the University-1652 dataset, there are 701 buildings, each containing one satellite view image and multiple drone view images. For the cross-view geo-localization task, the goal is to learn a mapping function that projects images from different platforms (e.g., drone and satellite) into a shared semantic space. In this space, the features of images representing the same location should be close to each other, while those from different locations should be well separated. This method enables effective matching of images from different views in the same location by leveraging feature similarity. It can address the challenges of significant viewpoint differences and supports accurate geo-localization.
@@ III-B Progressive Extension Embedding branch
Subsequently, adaptive average pooling (AvgPool) is applied to aggregate the multi-scale information, producing more discriminative semantic embeddings. Finally, the compact feature representation is processed through the batch normalization operation , followed by a dropout operation for regularization to aviod overfitting, and then propagated to the linear layer for classification. We outline these processing operations , , , as a lump
@@ III-C Global Extension Embedding Branch
The GEE branch adopts a global context aggregation strategy compared to the PEE branch, which progressively generates and enhances local embedded details. In the GEE branch, a mean pooling operation replaces the DEG module, processing the output features of the backbone to obtain a global feature representation. The global feature representation is then propagated to the DEC module for further refinement.
@@
where, and are height factor and width factor, respectively. This operation compresses the spatial dimensions, where . The global feature is then fed into the DEC module for further processing to obtain described as the DEC module III-B. In this way, the GEE branch can aggregate global information, generating comprehensive global feature embeddings that provide the DEC module with context-aware global features. Based on this global perspective, global and local consistency is ensured in the GEE branch.
@@ III-D Cross-Domain Enhanced Alignment branch
In cross-view geo-localization tasks, images taken from different platforms often exhibit significant geometric differences, resulting in substantial variations in feature distributions and posing challenges for cross-domain matching. Feature alignment has been proven to be effective in addressing this issue[31, 32, 33]. Inspired by [19], we designed a cross-domain enhanced alignment (CEA) branch that uses multi-level fusion and adaptive calibration strategies to dynamically adjust feature consistency within a shared latent space in high-dimensional embeddings. In addition, invariance constraints are incorporated to learn the inherent structural invariances of these features. As shown in Fig.2 after the initial feature extraction by the backbone, we obtain high-dimensional features from different perspectives (drone view and satellite view ). To efficiently utilize global information and facilitate cross-view feature alignment and matching, we apply a transformation to flatten the spatial dimensions into a single dimension , obtaining reshaped features , which serve as the input features for the CEA branch. Here, , . Break down the subsequent procedures into three modules, including domain enhancement (DE) module, feature adaptive temperature (FAT) module, and multi-scale feature fusion (MSF) module, which are described as follows.